the scores of adults on an IQ test are approximately normal with a mean 100 and standard deviation 15. Corinne scores 118 on such a test. She scores higher than what percent of adults?
Answer #1

Given le=100 =15 P(x < 118) = P(Z < 118-100) = P(Z < 1.20) - 0.8849 she scored tigher than 88.49% of adults